OpenStudy (anonymous):

indefinite integral of 1/xln(x^3) dx

OpenStudy (anonymous):

I have work so far but as usual it is probably wrong...

OpenStudy (lgbasallote):

kind of....

OpenStudy (lgbasallote):

here's a hint.. let \[u = \ln (x^3)\] so \[ du = \frac 3xdx\] agree?

OpenStudy (anonymous):

Yeah.. so one gets integral 1/u*(3/xdx)

OpenStudy (anonymous):

and then do you move the 3/x to the outside?

OpenStudy (lgbasallote):

hmm wait.. im getting confused with these x and u's combined

OpenStudy (lgbasallote):

\[du = \frac 3x dx\] so \[\frac{du}{3} = \frac{dx}{x}\] \[\int \frac{dx}{x \ln x^3} \implies \int \frac{du}{3u}\] right?

OpenStudy (anonymous):

Oh, okay... yeah

OpenStudy (lgbasallote):

so take out 1/3 out of the integral \[\implies \frac 13\int \frac{du}{u}\] now this is basic integration right?

OpenStudy (anonymous):

right because integral u/du is the same as integral 1/u du which is log u

OpenStudy (anonymous):

sub for u and multiply by 1/3?

OpenStudy (lgbasallote):

right.

OpenStudy (anonymous):

meant du/u* okay thanks

OpenStudy (lgbasallote):

welcome